OpenStudy (anonymous):

the 6th term of an arithmetic series is twice the 3rd term and the first term is 3. find the common difference and the 10th term. the 6th term of an arithmetic series is twice the 3rd term and the first term is 3. find the common difference and the 10th term. @Mathematics

OpenStudy (anonymous):

you should now be able 2 solve this one on your own..

OpenStudy (anonymous):

correctly said ap is the best thing in maths

OpenStudy (anonymous):

The common difference is 2 ?

OpenStudy (anonymous):

The 10th term is 2, right ?

OpenStudy (anonymous):

sorry i meant 21

OpenStudy (anonymous):

d=2 and a10 is 30

OpenStudy (anonymous):

if d=2 then a3=7....a6=13??

OpenStudy (anonymous):

d is 3 i am sorry

OpenStudy (anonymous):

my mistake

OpenStudy (anonymous):

could you explain how you got 3, i want to see where i got it wrong..

OpenStudy (anonymous):

a6=a+(n-1)d a3=a+(n-!)d a+5d=2a+4d

OpenStudy (anonymous):

d=a

OpenStudy (anonymous):

and so its a=3

OpenStudy (anonymous):

oh ! got it ! thanks ! :)